# Modelling the Spread of New Information on Social Networks

### v3

@@ 3.3.1 User Profile Features (U-P)
These features include the number of followers/followees of a user [1, 40, 47, 2], the number of lists that a user is included in [16, 15], the number of messages a user has posted since its registration [16, 48, 40], whether a user is verified [21, 15], whether a URL is provided in a user’s profile [39], a user’s indegree [13] and LeaderRank [13], and whether a user is following another user [15].
@@ 3.3.2 User Historical Action Features (U-HA)
These features include, for example, the percentage of a user’s historical messages that are reposts [21, 13], the average time interval between a user’s posts [21], and the ratio of a sender’s historical messages that have mentioned a recipient [16, 15].
3.3.3 User Historical Message Features (U-HM)
These features include the aggregated message features (M) of a sender and those of a recipient, as well as an additional feature measuring topic similarity between the two users [13, 20].
@@ 3.4 Positive and Negative Instances
In this study, we define a positive instance as a reposting event $\{M(h),U_{S},U_{R},t\}$, where a recipient $U_{R}$, at time $t$, reposted a message $M(h)$ with a hashtag $h$, and the message was created by a sender $U_{S}$ within the 24 hours before $t$.
Correspondingly, a negative instance is $\{M^{*}(h),U^{*}_{S},U_{R},t\}$, where the same recipient $U_{R}$ did not repost another message $M^{*}(h)$ containing the same hashtag $h$, and the message $M^{*}(h)$ was created by sender $U^{*}_{S}$ (which can be the same as $U_{S}$) within the 24 hours before the same $t$. The messages in the positive and negative instances contained the same hashtag and were created in the same 24 hours, so that it is likely that $M^{*}(h)$ was also exposed to the recipient and therefore make the evaluation more relevant.
@@
  • 1:1 dataset: For each positive instance, one negative instance is selected from our X data so that the message and the sender in the negative instance are most similar to those in the positive instance, i.e. the cosine distance between the feature vectors (containing all features of the message and the sender) of the positive instance and the chosen negative instance is the smallest among all negative instances.
  • 1:5 dataset: For each positive instance, the five most similar negative instances are selected, as defined above.
@@
Note that in all experiments in this study, if data is randomly split into training (and validation) and testing datasets, we remove any instance from the training dataset if the instance’s sender-recipient pair appears in the testing dataset. Such that, instances in the training and those in testing datasets always have different sender-recipient pairs. This is to prevent information leakage and ensure a fair and rigorous evaluation.
3.5 Prediction Models
3.5.1 Decision Tree (DT) Models
The DT models remain a competitive choice for reposting prediction [13, 15]. In this study, we will examine how the performance of DT models are affected when different types of features are used. We use the XGBoost decision tree model. We searched for the optimal hyper-parameters for each data split of each dataset. The most frequently used hyper-parameters include the maximum depth of 8, the learning rate of 0.3, the estimators of 100, the minimum child weight of 1, the subsample ratio of 1.0, and the positive class weight of 1.0. For more details, please refer to the Appendix.
3.5.2 Neural Network (NN) Models
@@
As shown in Figure 1, the NN models used in this study are extended from the SUA-ACNN model [14]  [^1] The SUA-ACNN model only used textual data, but it used textual data from not only the message being considered for reposting, Data(M), but also historical messages of the sender and the recipient of the reposting, Data(U-HM). In our study, we separate message data from user-related data. by adding a component of Multi-Layer Perceptron (MLP)  [^2] The MLP component has three fully connected layers with 128, 128 and 64 units, respectively. The activation functions are set to ReLU. to incorporate additional user-related data, i.e. Data(U-P) and Data(U-HA). This allowed us to obtain a series of models using different types of data, including the NN-U model using the three user-related data, the NN-M model using the message data only, and the NN-ALL model using all types of data. NN models are trained with the Adam optimizer, the categorical cross-entropy loss, the learning rate of $0.001$, the batch size of $40$, and early stopping on validation loss.
3.5.3 BERT
We also study the BERT model [49], which is a widely studied Large Language Model (LLM). The model takes the message data as input and generates embeddings that represent the semantic meanings of the messages. The model then produced the prediction based on the embeddings. We configure BERT-base with the AdamW optimizer, the learning rate of 2e52e-5, the batch size of $32$, and the dropout rate of $0.1$, with early stopping on validation loss. Cross-entropy loss was used as the objective function.
3.6 Experiments
In this study, we conducted three experiments for reposting prediction using the above models. We examined whether the performance of reposting prediction can be improved by supplementing or replacing message-related data with user-related data. In addition, we evaluated the models for both in-distribution and out-of-distribution predictions.
3.6.1 F1 Score
The formula of F1 Score is shown below, where $TP$, FPFP, and $FN$ represent true positives, false positives, and false negatives, respectively.
$F1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}$ (2)
3.6.2 Mean and Standard Deviation
For each prediction setting, we obtain ten F1 Scores. We calculate the mean $\boldsymbol{\mu}$ and standard deviation $\boldsymbol{\sigma}$ as defined below, where $j$ indexes the ten F1 Scores.
$\mu=\frac{1}{10}\sum_{j=1}^{10}F1_{j}$ (3)
$\sigma=\sqrt{\frac{1}{10}\sum_{j=1}^{10}(F1_{j}-\mu)^{2}}$ (4)
3.6.3 Overall Mean and Overall Standard Deviation
When predicting for an individual hashtag (in Experiments II and III), we obtain 10 F1 Scores and calculate the mean and the standard deviation as above, and then we further compute the overall mean $\boldsymbol{\bar{\mu}}$ and overall standard deviation $\boldsymbol{\bar{\sigma}}$ by treating results from the 14 hashtags as equally weighted components of a mixture distribution, defined as followings where $i$ indexes the 14 hashtags.
$\bar{\mu}=\frac{1}{14}\sum_{i=1}^{14}\mu_{i}$ (5)
4 Experiment I: In-Distribution Prediction For Mixed Hashtags
In this experiment, we mix data of all hashtags, and then performed Monte Carlo cross-validation 10 times, each randomly splitting the data 63/7/30 to the training, validation and test sets.
Table 2: In-Distribution Prediction for Mixed Hashtags
@@ 4 Experiment I: In-Distribution Prediction For Mixed Hashtags
Feature(ALL2) Feature(U3) Feature(M) Data(All) Data(U) Data(M) Data(M)
Dataset 1 F1 Score ($\mu$ $\pm$ $\sigma$4
@@
1 Ratio of positive and negative instances. 2 ALL include both user-related features U and message-related features M. 3 U includes U-P, U-HA and U-HM. 4 The mean $\mu$ and standard deviation $\sigma$ are computed as in Eq. (3)-(4).
4.1 Results
As shown in Table 2, the DT-ALL model using all types of features (303 in total) consistently shows significantly improved performance than other models using fewer features across datasets with different ratios of positive and negative instances. Among models with reduced feature sets, the DT-U model (utilizing user-related features) outperforms the DT-M model (utilizing message-related features).
The NN-ALL model utilizing all types of data produces the strongest performance among NN models. The NN-M model and the BERT model both utilize only message-related data and they achieve similar performance. Notably, these two models are outperformed by the NN-U model which used only user-related data.
These results consistently indicate that the inclusion of user-related data is valuable for reposting prediction.
As expected, the performance of all models degrades as the proportion of negative instances increases in the dataset, which aligns with established understanding of classification challenges in imbalanced datasets. In the following of this paper, all results are based on the 1:5 dataset.
4.2 Comparison with Prior Works
In addition to the above results, we also conducted Experiment I using the decision tree model, TORS, in [13], which uses 37 features primarily derived from textual data of messages, and the neural network model, SUA-ACNN, in [14], which relies solely on textual data. On the 1:5 dataset, they achieved F1 scores of 0.737 and 0.739, respectively, which as expected are close to but lower than the DT-M model and the NN-M model.
Table 3: Top 20 Most Important Features based on Decision Tree Prediction.
@@ 4.2 Comparison with Prior Works
Number of features in each type: U-HA: 10; U-HM: 4; U-P: 3; M: 3.
1 M_TopicLDA7 represents topics related to energy crisis, Brexit, and health issues. See full details of features in the Appendix.
4.3 Analysis of Feature Importance
To further analyse the relative importance of the 303 features, we obtained their importance values as a by-product of XGBoost decision tree prediction. The feature importance value represents the relative contribution of a feature to the decision tree model’s predictive performance. Specifically, it is computed as the reduction in the loss function contributed by each feature when splitting nodes, with larger reductions indicating higher importance. The feature importance values are normalised such that the sum for all features equals to one.
We obtained the feature importance values for all 303 features by training the DT-ALL model using all (instead of 70%) data in the 1:5 dataset. Table 3 shows the top 20 most important features. It is notable that the majority of top 20 features are user-related features, including 10 user historical action (U-HA) features, 4 user historical message (U-HM) features, and 3 user profile (U-P) features. The most important message (M) feature is merely ranked at the 6th position. This highlights the importance of user-related features for reposting predictions.
5 Experiment II: In-Distribution Prediction For Individual Hashtags
In this experiment, we conduct the in-distribution predictions for each of the 14 individual hashtags. We performed the same Monte Carlo cross-validation but on data of a single hashtag.
Table 5 presents the in-distribution prediction results for each individual hashtag. The results are broadly consistent with those observed from the mixed-hashtag setting, with minor variations across individual hashtags. Specifically, the overall ranking of models remains the same: for decision trees, DT-ALL outperforms DT-U, which in turn outperforms DT-M; for neural network models, NN-ALL achieves the highest performance, followed by NN-U, with NN-M and BERT producing similar lowest performance. These findings indicate that the relative effectiveness of user-related and message-related data is stable across different in-distribution conditions.
@@ 5 Experiment II: In-Distribution Prediction For Individual Hashtags
Note that, for rigorous evaluation, we have also conducted additional temporal experiments where data of each individual hashtag are divided into time windows according to the timestamp of repost events (see additional temporal experiments in Appendix). The results and conclusions remain the same: models using user-related data consistently outperform those using message-related data.
Table 4: In-Distribution Prediction for Individual Hashtags on the 1:5 Dataset
In-Distribution For a given hashtag, training on 70% randomly selected data, test on remaining 30% data.
Decision Tree (DT) model Neural Network (NN) model LLM model
@@
The mean, $\mu$, and standard deviation, $\sigma$, of F1 score for each hashtag are computed using Eq. (3)-(4). The overall mean, $\bar{\mu}$, and overall standard deviation, $\bar{\sigma}$, across 14 hashtags are computed using Eq. (5)-(6). Same in Table 5.
Table 5: Out-of-Distribution Prediction on the 1:5 Dataset
Out-of-distribution For a given hashtag, training on data of other hashtags, test on data of given hashtag.
@@
The expected F1 score for random guessing is 0.167, where positive and negative labels are randomly assigned to data at a ratio of 1:5.
6 Experiment III: Out-of-Distribution Prediction
We conducted out-of-distribution prediction for each of the 14 hashtags in our X data. For each given hashtag, we trained a model using data of the other 13 hashtags, and then tested the model with data of the given hashtag (the new, unseen topic). This split ensures an out-of-distribution condition, where the test messages come from a hashtag (topic) unseen during training  [^3] In our X data, we removed the small amount (less than 0.1%) of sender–recipient pairs that are overlapping across different hashtags. This ensures that sender–recipient pairs in the testing data are also unseen during training..
To obtain a stable and reliable estimate of model performance, for each given hashtag, we randomly divided the training data of the other 13 hashtags into 10 subsets, left one subset out for validation and used the other 9 subsets to train the model, and obtain prediction results. Then we repeated the process 10 times by leaving a different subset out each time. We evaluated the 10 trained models against the same testing data of the given hashtag under study, and then obtained the mean and standard deviation from ten F1 scores.
6.1 Results
Table 5 shows the out-of-distribution prediction by DT, NN and BERT models for each of the 14 hashtags on the 1:5 dataset. When based only on message-related data, all models show notably poor performance. Their overall F1 scores, $\bar{\mu}$, across 14 hashtags are as low as 0.117, 0.108 and 0.091, respectively, which are close to the expected F1 score of random guessing (0.167, see notes in Table 5). By comparison, the DT-ALL model and the NN-ALL model, using all types of data, perform significantly better.
Remarkably, the DT-U model and NN-U models, using only the user-related data, produce performance as good as, or even better than the DT-ALL and NN-ALL using all data. This suggests that user-related data are crucial for out-of-distribution prediction, where adding message-related data would have minimum benefit, or even negative impact, on performance.
The low performance of the NN-M model and the BERT model are unlikely due to any limitation in their architecture. Their underperformance is a result of exclusive reliance on message data. These two models learn from textual content of messages and therefore are bounded by the vocabularies present in the training data. This constrains their ability to generalise to messages of a new topic that are absent in the training data.
6.2 Discussion
6.2.1 In-Distribution vs. Out-of-Distribution

### v4

@@ 3.3.1 User Profile Features (U-P)
These features include the number of followers/followees of a user [40, 37, 64, 12], the number of lists that a user is included in [34, 53], the number of posts a user has created since its registration [34, 24, 37], whether a user is verified [19, 53], whether a URL is provided in a user’s profile [6], a user’s indegree [65] and LeaderRank [65], and whether a user is following another user [53].
@@ 3.3.2 User Historical Action Features (U-HA)
These features include, for example, the percentage of a user’s historical posts that are reposts [19, 65], the average time interval between a user’s posts [19], and the ratio of a sender’s historical posts that have mentioned a recipient [34, 53].
3.3.3 User Historical Post Features (U-HM)
These features include the aggregated post features (M) of a sender and those of a recipient, as well as an additional feature measuring topic similarity between the two users [65, 42].
@@ 3.4 Positive and Negative Instances
In this study, we define a positive instance as a reposting event $(M_{h},U_{S,t_{s}},U_{R,t_{r}})$, where the (original) post $M_{h}$ with a hashtag $h$ was created by a sender $U_{S}$ at $t_{s}$, and was reposted by a recipient $U_{R}$ at time $t_{r}$ (within 24 hours since $t_{s}$). Correspondingly, a negative instance is $(M^{*}_{h},U^{*}_{S,t_{s}},U_{R,t_{r}})$, where the same recipient $U_{R}$ did not repost another post $M^{*}_{h}$ with the same hashtag $h$ created by sender $U^{*}_{S}$ (where it is possible $U^{*}_{S}=U_{S}$). Here we consider negative instances that contained the same hashtag in the positive case and was created in the same 24-hour window, such that it is likely that $M^{*}_{h}$ was also exposed to the recipient and therefore makes the evaluation more relevant.
@@
  • 1:1 dataset: For each positive instance, one negative instance is selected from our X data so that the post and the sender in the negative instance are most similar to those in the positive instance, i.e., the cosine distance between the feature vectors (containing all features of the post and the sender) of the positive instance and the chosen negative instance is the smallest among all negative instances.
  • 1:5 dataset: For each positive instance, the five most similar negative instances are selected, as defined above. This ratio is typically used in such experiments, such as in [60].
@@
Note that in all experiments in this study, if data is randomly split into training (and validation) and testing datasets, we remove any instance from the training dataset if the instance’s sender-recipient pair appears in the testing dataset. Such that, instances in the training and those in the testing datasets always have different sender-recipient pairs. This is to prevent information leakage and ensure a fair and rigorous evaluation.
4 Models for Reposting Prediction
In this study, we aim to examine the different contributions of post data and user data to the reposting prediction under both in-distribution and out-of-distribution settings. To focus on the aim, we studied three standard classification models to avoid complex tuning and optimisation associated with state-of-the-art models.
4.1 Decision Tree (DT) Models
The DT models remain a competitive choice for reposting prediction [65, 53]. In this study, we will examine how the performance of DT models is affected when different types of features are used. We use the XGBoost decision tree model. We searched for the optimal hyper-parameters for each data split of each dataset. The most frequently used hyper-parameters include the maximum depth of 8, the learning rate of 0.3, the estimators of 100, the minimum child weight of 1, the subsample ratio of 1.0, and the positive class weight of 1.0. For more details, please refer to the Appendix.
4.2 Neural Network (NN) Models
@@
As shown in Figure 1, the NN models used in this study are extended from the SUA-ACNN model111 The SUA-ACNN model only used textual data, but it used textual data from not only the post being considered for reposting, Data(M), but also historical posts of the sender and the recipient of the reposting, Data(U-HM). In our study, we separate post data from user data.[60] by adding a component of Multi-Layer Perceptron (MLP)222 The MLP component has three dense layers with 128, 128 and 64 units, respectively. The activation function is set to ReLU. to incorporate additional user data, i.e., Data(U-P) and Data(U-HA). This allowed us to obtain a series of models using different types of data, including the NN-U model using the user data, the NN-M model using the post data only, and the NN-ALL model using all types of data. We set the hyperparameters as in [60]. NN models are trained with the Adam optimizer, the categorical cross-entropy loss, the learning rate of $0.001$, the batch size of $40$, and early stopping on validation loss.
4.3 BERT
We also study the BERT model [14], which is a widely studied Large Language Model (LLM). The model takes the post data as input and generates embeddings that represent the semantic meanings of the posts. The model then produced the prediction based on the embeddings. We used the recommended fine-tuning settings from [14]. We configure BERT-base with the AdamW optimizer, the learning rate of 2e52e-5, the batch size of $32$, and the dropout rate of $0.1$, with early stopping on validation loss. Cross-entropy loss was used as the objective function.
5 Experiments and Results
In this study, we conducted three experiments for reposting prediction using the above models. We examined whether the performance of reposting prediction, in terms of F1 score, can be improved by supplementing or replacing post data with user data. Firstly, we evaluated the models for in-distribution prediction on data of mixed hashtags; secondly, we compared the in-distribution prediction for individual hashtags; and thirdly, we conducted the out-of-distribution predictions.
See Appendix for definitions of the mean $\mu_{i}$, the standard deviation $\sigma_{i}$, the overall mean $\bar{\mu}$, and the overall standard deviation $\bar{\sigma}$.
5.1 Experiment I: In-Distribution Prediction For Mixed Hashtags
In this experiment, we mix data of all hashtags, and then performed Monte Carlo cross-validation 10 times, each randomly splitting the data 63/7/30 to the training, validation, and test sets.
@@ 5.1 Experiment I: In-Distribution Prediction For Mixed Hashtags
Dataset F1 Score ($\mu$ $\pm$ $\sigma$)
@@
Table 2: In-distribution prediction for mixed hashtags. Datasets are named by the ratio of positive and negative instances. Feature(ALL) includes both post-related features, Feature(M), and user-related features, Feature(U) which includes U-P, U-HA, and U-HM.
As shown in Table 2, the DT-ALL model using all types of features (303 in total) consistently shows significantly improved performance than other models using fewer features across datasets with different ratios of positive and negative instances. Among models with reduced feature sets, the DT-U model (utilizing user-related features) outperforms the DT-M model (utilizing post-related features).
The NN-ALL model utilizing all types of data produces the strongest performance among NN models. The NN-M model and the BERT model both utilize only post data and they achieve similar performance. Notably, these two models are outperformed by the NN-U model which used only user data.
These results consistently indicate that the inclusion of user data is valuable for reposting prediction.
As expected, the performance of all models degrades as the proportion of negative instances increases in the dataset, which aligns with the established understanding of classification challenges in imbalanced datasets. In the following of this paper, all results are based on the 1:5 dataset.
5.1.1 Comparison with Prior Works
In addition to the above results, we also conducted Experiment I using the decision tree model, TORS, in [65], which uses 37 features primarily derived from post data, and the neural network model, SUA-ACNN, in [60], which relies solely on post data. They achieved F1 scores of 0.737 and 0.739, respectively, which as expected are close to but lower than the DT-M model and the NN-M model.
@@ 5.1.1 Comparison with Prior Works
# of features in each type: U-HA(10), U-HM(4), U-P(3), M(3).
Table 3: Top 20 most important features based on Decision Tree prediction. 1M_TopicLDA7 represents topics related to energy crisis, Brexit, and health issues. See full details of features in the Appendix.
5.1.2 Feature Importance
The feature importance value represents the relative contribution of a feature to the decision tree model’s predictive performance. Specifically, it is computed as the reduction in the loss function contributed by each feature when splitting nodes, with larger reductions indicating higher importance.
We obtained the feature importance values for all 303 features by training the DT-ALL model using all (instead of 70%) data in the 1:5 dataset. The feature importance values are normalised such that the sum for all features equals one. Our analysis showed that less than 1% of feature pairs has a Pearson coefficient value higher than 0.7 [16], which means most features contribute to the prediction.
Table 3 shows the top 20 most important features. It is notable that the majority of the top 20 features are user-related features, including 10 user historical action (U-HA) features, 4 user historical post (U-HM) features, and 3 user profile (U-P) features. The most important post (M) feature is merely ranked at the 6th position. This highlights the importance of user-related features for reposting predictions.
5.2 Experiment II: In-Distribution Prediction For Individual Hashtags
In this experiment, we conduct the in-distribution predictions for each of the 14 hashtags. We performed the same Monte Carlo cross-validation but on data of an individual hashtag.
Table 5 presents the in-distribution prediction results for each individual hashtag. The results are broadly consistent with those observed from the mixed-hashtag setting, with minor variations across individual hashtags. Specifically, the overall ranking of models remains the same: for decision trees, DT-ALL outperforms DT-U, which in turn outperforms DT-M; for neural network models, NN-ALL achieves the highest performance, followed by NN-U, with NN-M and BERT producing similar lowest performance. Both paired t-tests and Wilcoxon signed-rank tests confirm that the above performance rankings are statistically significant (with all values of p << 0.05). These findings indicate that the relative effectiveness of user and post data is stable across different in-distribution conditions.
@@ 5.2 Experiment II: In-Distribution Prediction For Individual Hashtags
Note that, for rigorous evaluation, we have also conducted additional temporal experiments where data of each individual hashtag are divided into time windows according to the timestamp of repost events (see additional temporal experiments in Appendix). The results and conclusions remain the same: models using user data consistently outperform those using post data.
In-Distribution prediction For a given hashtag, training on 70% randomly selected data, test on remaining 30% data.
Decision Tree (DT) models Neural Network (NN) model LLM model
@@
Table 4: In-distribution prediction for individual hashtags on the 1:5 dataset. See Appendix for the calcuation of $\mu$, $\sigma$, the overall mean $\bar{\mu}$, and the overall standard deviation $\bar{\sigma}$.
Out-of-distribution prediction For a given hashtag, training on data of the other 13 hashtags, test on the given hashtag.
@@
Table 5: Out-of-distribution prediction for individual hashtags on the 1:5 dataset. The expected F1 score for random guessing is 0.167, where positive and negative labels are randomly assigned at a ratio of 1:5.
5.3 Experiment III: Out-of-Distribution Prediction
We conducted out-of-distribution prediction for each of the 14 hashtags in our X data. For each given hashtag, we trained a model using data of the other 13 hashtags, and then tested the model with data of the given hashtag (the new, unseen topic). This split ensures an out-of-distribution condition, where the test posts come from a hashtag (topic) unseen during training 333 In our X data, we removed the small amount (less than 0.1%) of sender–recipient pairs that are overlapping across different hashtags. This ensures that sender–recipient pairs in the testing data are also unseen during training..
To obtain a stable and reliable estimate of model performance, for each given hashtag, we randomly divided the training data of the other 13 hashtags into 10 subsets, left one subset out for validation and used the other 9 subsets to train the model, and obtain prediction results. Then we repeated the process 10 times by leaving a different subset out each time. We evaluated the 10 trained models against the same test data of the given hashtag under study, and then obtained the mean and standard deviation of F1 scores.
Table 5 shows the out-of-distribution prediction by DT, NN and BERT models for each of the 14 hashtags on the 1:5 dataset. When based only on post data, all models show notably poor performance. Their overall F1 scores, $\bar{\mu}$, across 14 hashtags are as low as 0.117, 0.108, and 0.091, respectively, which are close to the expected F1 score of random guessing (0.167, see notes in Table 5). By comparison, the DT-ALL model and the NN-ALL model, using all types of data, perform significantly better.
Remarkably, the DT-U model and NN-U models, using only the user-related data, produce performance as good as, or even better than the DT-ALL and NN-ALL using all data. This suggests that user-related data are crucial for out-of-distribution prediction, where adding post-related data would have minimum benefit, or even negative impact, on performance. Both paired t-tests and Wilcoxon signed-rank tests confirm the performance rankings of DT-U >> DT-M, DT-ALL >> DT-M, and NN-U >> NN-ALL >> NN-M are statistically significant (with all values of p << 0.05).
The low performance of the NN-M model and the BERT model is unlikely due to any limitation in their architecture. Their underperformance is a result of exclusive reliance on post data. These two models learn from textual content of posts and therefore are bounded by the vocabularies present in the training data. This constrains their ability to generalise to posts of a new topic that are absent in the training data.
6 Discussion
6.1 In-Distribution vs. Out-of-Distribution